# Interplanetary magnetic correlation and low-frequency spectrum over many solar rotations

### v1

@@ Abstract
{CJK*}
UTF8gbsn
1 Introduction
Statistical methods have long been employed to study variability of heliospheric plasma and magnetic fields. Correlation and spectral analysis have been prominent in this regard, particularly in characterizing properties of turbulence (Bruno & Carbone, 2013). Aided by the Taylor hypothesis that approximately relates temporal fluctuations at a fixed point to spatial fluctuations advected by the flow (Taylor, 1938), spectral analysis enables study of turbulence properties. Prominent among these is the familiar power-law inertial range typically interpreted as evidence of a local Kolmogorov-like energy cascade (Coleman, 1968). At lower frequencies, below the measured correlation timescale, one might expect a flat spectrum indicative of completely uncorrelated signals. Yet, this is not often seen in the interplanetary magnetic field. Instead, for very low frequencies (typically below $10^{-3}\,\mathrm{Hz}$), a different feature called $1/f$ noise emerges, suggestive of persistent correlation. Also known as “flicker noise”, the $1/f$ spectrum has the unique property of distributing equal power across logarithmic frequency intervals, i.e., constant power per octave. While there exist generic pathways to generating the $1/f$ signal in a variety of physical systems (Machlup, 1981; Dutta & Horn, 1981; Montroll & Shlesinger, 1982; Bak et al., 1987), its origin in the heliospheric plasma remains a topic of debate (see, e.g., Matthaeus & Goldstein (1986); Velli et al. (1989); Verdini et al. (2012); Matteini et al. (2018); Chandran (2018); Magyar & Van Doorsselaere (2022) for various proposed generation mechanisms; and Davis et al. (2023); Huang et al. (2023); Pradata et al. (2025) for some recent observations). See Wang et al. (2024a) for a recent review.
@@ 1 Introduction
In the present paper, we examine the context of the $1/f$ signal, emphasizing its relationship to the turbulence inertial range and also to signals at very low frequencies, in particular those having periods related to solar rotation. We show in Section 3 the presence of a secondary 13.5-day periodicity superposed on the dominant 27-day solar rotation periodicity, which becomes particularly prominent during the solar minimum phase of the analyzed 23rd solar cycle. Periodic signals on day-long timescales in the solar wind, especially those associated with solar rotation or linked to chromospheric structures, have been extensively studied for their role in shaping solar-terrestrial interactions (see, e.g., Xie et al., 2017).
@@
This paper is organized as follows: in Section 2, we describe the data employed, the correlation analysis, and the power-spectral analysis. In Section 3, we compare results from correlation and spectral analysis over different solar wind categories, including the slow and fast solar wind, as well as during solar maximum and minimum. In Section 4, we discuss our results in the context of the discussion posed above.
2 Data and Analysis Procedure
@@
[…]
@@
[…]
@@
[…]
is evaluated via the discrete Fourier transform of the bounded, discretely sampled correlation function $R(\tau)$. To prevent $R(\tau)$ from being artificially treated as periodic over the finite interval $[-T,T]$, we first apply a 10% cosine-taper window (Matthaeus & Goldstein, 1982a), in which the final 10% of the correlation values are multiplied by the factor
@@
3 Results
3.1 Magnetic correlation
Most past work on the magnetic autocorrelation, as reviewed in Section 1, has reported signals at 27 days and a few harmonics in the spectrum, but only a few have analyzed the correlation functions using data records long enough to probe time lags of many months (Gosling et al., 1976; Mursula & Zieger, 1996; Xie et al., 2017). Here, our emphasis is on the nature of very low frequency signals, and accordingly, our methodology employs correlation analysis with lags up to several years. We present the analysis in several stages.
@@ 3.1 Magnetic correlation
The correlation functions computed separately for the solar maximum (1999-2002) and minimum (2006-2009) phases are shown in Fig. 3. During solar minimum, a clear 13.5-day periodicity, corresponding to half the synodic solar rotation period, emerges alongside the primary 27-day signal. This feature has been reported previously as described in Section 1. In contrast, the correlations during solar maximum display no distinct secondary structures. However, they begin to shift out of phase with the 27-day periodicity after approximately five solar rotations, instead exhibiting a slightly shorter period closer to around 26 days. This alludes to the possible existence of a range of periodicities, which we will examine in greater depth below.
@@
3.2 Magnetic spectrum
The panels of Fig. 6 illustrate power spectral densities of magnetic field components (left column) and the normalized trace spectrum (right column) computed from the autocorrelations. The rows show spectra from each of the five analyzed wind categories: (1) the full 12-year dataset, (2) slow wind, (3) fast wind, (4) the solar 1999-2002 maximum, and (5) the 2006-2009 solar minimum. Each panel is divided into the corresponding spectral density itself (top half), and the associated coarse-grained (integrated in frequency bins) spectra (bottom half).
@@ 3.2 Magnetic spectrum
To examine spectral scaling across different scales and to probe the extent of the $1/f$ range, we estimate the integrated spectra from the spectral densities using the method outlined in Section 2. The integrated spectra are plotted in the bottom panels of Fig. 6. The inertial range turbulence spectrum with a $f^{-5/3}$ power-law corresponds to a $f^{-2/3}$ scaling in the integrated spectrum, whereas the $1/f$ region appears flat, reflecting its characteristic property of equal power per octave (Machlup, 1981).
Across all five cases analyzed - namely the full 12-year dataset, slow wind, fast wind, solar minimum, and solar maximum - the radial and tangential magnetic field exhibit generally similar spectral behavior over the full range of frequencies studied (approximately $10^{-8}\,\mathrm{Hz}$ to $3\times 10^{-3}\,\mathrm{Hz}$, corresponding to timescales from around $3300\,\mathrm{seconds}$ to $3\,\mathrm{years}$). One exception is in the fast wind between $1.3\times 10^{-6}\,\mathrm{Hz}$ (the second superharmonic of solar rotation) and $10^{-5}\,\mathrm{Hz}$, where a slight deviation occurs. In general, the spectral indices for the radial and tangential components hover around $-1$ or marginally steeper immediately above the solar rotation frequency for at least a decade, before transitioning into the $f^{-5/3}$ inertial range. The normal component, by contrast, tends to remain flat - indicative of uncorrelated signals - before directly transitioning into the $f^{-5/3}$ regime, bypassing a clearly defined $1/f$ range. Therefore, the $1/f$ behavior observed in the trace spectrum for all five cases likely arises from the collective contribution of all three magnetic field components.
@@
3.3 Density correlation and spectrum
@@
4 Conclusion and Discussion
Our purpose has been to study in some detail the behavior of long-timescale, very-low-frequency interplanetary turbulence. A particular emphasis has been the connection between the fluctuation properties from the nominal correlation scale ($\sim$ one hour) to the nominal solar rotation scale ($\sim$ 27 days). To this end, we compute two-time correlation functions and spectra of the interplanetary magnetic field and proton density using a 12-year sample of ACE data. The focus has not been on structure in wavenumber, but rather structure in frequency, so we have not invoked the Taylor frozen-in hypothesis generally used to justify conversion to wavenumber (Jokipii, 1973). Working in frequency space, we do not attempt refined commentary on the spatial structure of the inertial range, usually expected to have a wavenumber dependence close to a Kolmogorov-like spectral index of $-5/3$. Nevertheless, we can recognize the inertial range form as near $f^{-5/3}$, in essentially all wind categories but the fast solar wind (see below). We also note that toward low frequencies at around $10^{-5}$ to $10^{-4}\,\mathrm{Hz}$, the spectral form gradually transitions to the $1/f$ regime. This extended transition region is consistent with the use of a long data record, along with the established large-variance log-normal distribution of interplanetary magnetic field correlation times (Ruiz et al., 2014; Wang et al., 2024a) and the nature of the superposition principle (Machlup, 1981; Montroll & Shlesinger, 1982) that provides a generic pathway to the generation of $1/f$ signal.
In the fast solar wind ($V_{\mathrm{SW}}\geq 500\,\mathrm{km/s}$), the inertial range spectral index – typically observed at frequencies above around $3\times 10^{-4}\,\mathrm{Hz}$ (if considering the lower bound of the 1 au correlation time distribution at around 1 hour) – deviates from $-5/3$ and flattens. This trend is evident across all three magnetic field components. Evidence of spectral flattening in fast wind has been presented by Borovsky (2012) employing ACE intervals categorized by wind speed and wind origin (e.g. coronal-hole, non-coronal-hole, and ejecta based on ionized oxygen ratio). Magnetic field spectral index is systematically less negative in fast wind and in wind of coronal-hole origin. Borovsky (2012) also identifies intervals of sustained shallow spectra reaching a spectral index of -1.34, which tend to have substantially enhanced velocity and magnetic field fluctuations, Alfvén ratios, and outward-to-inward Elsässer ratios. Teodorescu et al. (2015) using measurements at 0.72 au also find spectral flattening in fast wind, and attributes the observed slope tending toward $-3/2$ to strong Alfvénicity of the flow and limited turbulence development due to fast wind’s rapid transit. More recently, Sioulas et al. (2023) reports spectral indices near $-3/2$ independent of plasma speed closer to the Sun using Parker Solar Probe data.
@@ 4 Conclusion and Discussion
Appendix A Statistics on magnetic fluctuation variance
We present additional statistics on magnetic field fluctuation variance with dependence on solar wind speed and on solar cycle phase. As described in Section 2, the time series of magnetic component fluctuations $b_{i}$ are computed using a 6-hour sliding window average. Then instances of the average $\langle b^{2}\rangle=\langle b_{\mathrm{R}}^{2}+b_{\mathrm{T}}^{2}+b_{\mathrm{N}}^{2}\rangle$ are evaluated over consecutive 6-hour intervals. Fig. 10 shows the distribution $\langle b^{2}\rangle$ during the entire 23rd solar cycle, the 1999-2002 solar maximum, 2006-2009 solar maximum, as well as for slow and fast wind during both solar maximum and minimum.
Note that the fluctuation variances $\langle b^{2}\rangle$ are log-normally distributed across all cases examined. The mean of $\langle b^{2}\rangle$ is consistently highest in fast wind and during solar maximum, and lowest in slow wind and during solar minimum. The standard deviation of $\langle b^{2}\rangle$ systematically follow the exact same trend. The increase in $\langle b^{2}\rangle$ and its standard deviation in the fast solar wind may reflect enhanced fluctuations in the magnetic field vector rather than in its magnitude (Bruno & Carbone, 2013). The increase observed during solar maximum is consistent with previous studies (Borovsky, 2012; Zhao et al., 2018), and likely results from the heightened solar activity and irregular solar wind variability (Bruno & Carbone, 2013; D’Amicis & Bruno, 2015).
Figure 10: Distributions of magnetic field fluctuation variance over solar minimum (2006-2009), maximum (1999-2002), slow wind, fast wind, and full 12-year data. Expectation value and standard deviation of each distribution are listed in legend. Solid curves represent best-fit log-normal distribution with parameters also listed in legend.

### v2

@@ Abstract
I Introduction
tatistical methods have long been employed to study variability of heliospheric plasma and magnetic fields. Correlation and spectral analysis have been prominent in this regard, particularly in characterizing properties of turbulence (Bruno & Carbone, 2013). Aided by the Taylor hypothesis that approximately relates temporal fluctuations at a fixed point to spatial fluctuations advected by the flow (Taylor, 1938), spectral analysis enables study of turbulence properties. Prominent among these is the familiar power-law inertial range typically interpreted as evidence of a local Kolmogorov-like energy cascade (Coleman, 1968). At lower frequencies, below the measured correlation timescale, one might expect a flat spectrum indicative of completely uncorrelated signals. Yet, this is not often seen in the interplanetary magnetic field. Instead, for very low frequencies (typically below $10^{-3}\,\mathrm{Hz}$), a different feature called $1/f$ noise emerges, suggestive of persistent correlation. Also known as “flicker noise”, the $1/f$ spectrum has the unique property of distributing equal power across logarithmic frequency intervals, i.e., constant power per octave. While there exist generic pathways to generating the $1/f$ signal in a variety of physical systems (Machlup, 1981; Dutta & Horn, 1981; Montroll & Shlesinger, 1982; Bak et al., 1987), its origin in the heliospheric plasma remains a topic of debate (see, e.g., Matthaeus & Goldstein (1986); Velli et al. (1989); Verdini et al. (2012); Matteini et al. (2018); Chandran (2018); Magyar & Van Doorsselaere (2022) for various proposed generation mechanisms; and Davis et al. (2023); Huang et al. (2023); Pradata et al. (2025) for some recent observations). See Wang et al. (2024a) for a recent review.
@@ I Introduction
In the present paper, we examine the context of the $1/f$ signal, emphasizing its relationship to the turbulence inertial range and also to signals at very low frequencies, in particular those having periods related to solar rotation. We show in Section III the presence of a secondary 13.5-day periodicity superposed on the dominant 27-day solar rotation periodicity, which becomes particularly prominent during the solar minimum phase of the analyzed 23rd solar cycle. Periodic signals on day-long timescales in the solar wind, especially those associated with solar rotation or linked to chromospheric structures, have been extensively studied for their role in shaping solar-terrestrial interactions (see, e.g., Xie et al., 2017).
@@
This paper is organized as follows: in Section II, we describe the data employed, the correlation analysis, and the power-spectral analysis. In Section III, we compare results from correlation and spectral analysis over different solar wind categories, including the slow and fast solar wind, as well as during solar maximum and minimum. In Section IV, we discuss our results in the context of the discussion posed above.
II Data and Analysis Procedure
@@
is expressed in terms of ensemble averaged quantities and $\tau$ represents the temporal lag. In practice we proceed using the Blackman–Tukey algorithm (Blackman & Tukey, 1958; Matthaeus & Goldstein, 1982a), in which $\langle\cdots\rangle$ is interpreted as averaging over the entire duration of the specified data (see Wang et al., 2024b, for more details). The temporal lag $\tau$ spans the range from zero to $T$, where $T$ is set to 0.3 times the data duration to ensure sufficient statistical weight on each correlation value.
@@
To identity the 4-year subsets of solar minimum and maximum phases, we consult the sunspot number record, the standard indicator of solar activity (Clette & Lefèvre, 2015). The chosen solar maximum subset spans years 1999 to 2002, and the solar minimum subset spans 2006 to 2009. To further characterize the variability of magnetic field data in these phases, we analyze the magnetic field fluctuation variance $\langle b^{2}\rangle$ over time, as shown in Fig. 1. Here, the component-wise magnetic field fluctuation $b_{i}$ is first calculated using a boxcar sliding window of 6 hours (see Matthaeus & Goldstein, 1982a, for details on this method), approximately several times the correlation times at $1\,\mathrm{au}$ (see, e.g., Ruiz et al., 2014; Cuesta et al., 2022). The fluctuation variance is given by $\langle b^{2}\rangle=\langle b_{\mathrm{R}}^{2}\rangle+\langle b_{\mathrm{T}}^{2}\rangle+\langle b_{\mathrm{N}}^{2}\rangle$, where $\langle\cdots\rangle$ now denotes averaging over each consecutive 6-hour interval. As expected from heightened solar activity, the fluctuation variance is evidently higher during the solar maximum than during solar minimum. More statistics of the magnetic field fluctuation variance are shown in the SI Appendix.
@@
$S(f)=\int_{-\infty}^{\infty}R(\tau)e^{-i2\pi f\tau}d\tau,$ (3)
is evaluated via the discrete Fourier transform of the bounded, discretely sampled correlation function $R(\tau)$. To prevent $R(\tau)$ from being artificially treated as periodic over the finite interval [T,T][-T,T], we first apply a 10% cosine-taper window (Matthaeus & Goldstein, 1982a), in which the final 10% of the correlation values are multiplied by the factor
@@
III Results
III.1 Magnetic correlation
Most past work on the magnetic autocorrelation, as reviewed in the Introduction, has reported signals at 27 days and a few harmonics in the spectrum, but only a few have analyzed the correlation functions using data records long enough to probe time lags of many months (Gosling et al., 1976; Mursula & Zieger, 1996; Xie et al., 2017). Here, our emphasis is on the nature of very low frequency signals, and accordingly, our methodology employs correlation analysis with lags up to several years. We present the analysis in several stages.
@@ III.1 Magnetic correlation
The correlation functions computed separately for the solar maximum (1999-2002) and minimum (2006-2009) phases are shown in Fig. 3. During solar minimum, a clear 13.5-day periodicity, corresponding to half the synodic solar rotation period, emerges alongside the primary 27-day signal. This feature has been reported previously as described in the Introduction. In contrast, the correlations during solar maximum display no distinct secondary structures. However, they begin to shift out of phase with the 27-day periodicity after approximately five solar rotations, instead exhibiting a slightly shorter period closer to around 26 days. This alludes to the possible existence of a range of periodicities, which we will examine in greater depth below.
@@
III.2 Magnetic spectrum
The panels of Fig. 6 illustrate power spectral densities of magnetic field components (left column) and the normalized trace spectrum (right column) computed from the autocorrelations. The rows show spectra from each of the five analyzed wind categories: (1) the full 12-year dataset, (2) slow wind, (3) fast wind, (4) the solar 1999-2002 maximum, and (5) the 2006-2009 solar minimum. Each panel is divided into the corresponding spectral density itself (top half), and the associated coarse-grained (integrated in frequency bins) spectra (bottom half). Compensated trace spectra $f\hat{S}$, plotted in both log-log and linear-log formats, are provided in the SI Appendix.
@@ III.2 Magnetic spectrum
To examine spectral scaling across different scales and to probe the extent of the $1/f$ range, we estimate the integrated spectra from the spectral densities using the method outlined in Section II. The integrated spectra are plotted in the bottom panels of Fig. 6. The inertial range turbulence spectrum with a $f^{-5/3}$ power-law corresponds to a $f^{-2/3}$ scaling in the integrated spectrum, whereas the $1/f$ region appears flat, reflecting its characteristic property of equal power per octave (Machlup, 1981).
Across all five cases analyzed - namely the full 12-year dataset, slow wind, fast wind, solar minimum, and solar maximum - the radial and tangential magnetic field exhibit generally similar spectral behavior over the full range of frequencies studied (approximately $10^{-8}\,\mathrm{Hz}$ to $3\times 10^{-3}\,\mathrm{Hz}$, corresponding to timescales from around $3300\,\mathrm{seconds}$ to $3\,\mathrm{years}$). One exception is in the fast wind between $1.3\times 10^{-6}\,\mathrm{Hz}$ (the second superharmonic of solar rotation) and $10^{-5}\,\mathrm{Hz}$, where a slight deviation occurs. In general, the spectral indices for the radial and tangential components hover around 1-1 or marginally steeper immediately above the solar rotation frequency for at least a decade, before transitioning into the $f^{-5/3}$ inertial range. The normal component, by contrast, tends to remain flat - indicative of uncorrelated signals - before directly transitioning into the $f^{-5/3}$ regime, bypassing a clearly defined $1/f$ range. Therefore, the $1/f$ behavior observed in the trace spectrum for all five cases likely arises from the collective contribution of all three magnetic field components.
@@
III.3 Density correlation and spectrum
@@
IV Conclusion and Discussion
Our purpose has been to study in some detail the behavior of long-timescale, very-low-frequency interplanetary turbulence. A particular emphasis has been the connection between the fluctuation properties from the nominal correlation scale ($\sim$ one hour) to the nominal solar rotation scale ($\sim$ 27 days). To this end, we compute two-time correlation functions and spectra of the interplanetary magnetic field and proton density using a 12-year sample of ACE data. The focus has not been on structure in wavenumber, but rather structure in frequency, so we have not invoked the Taylor frozen-in hypothesis generally used to justify conversion to wavenumber (Jokipii, 1973). Working in frequency space, we do not attempt refined commentary on the spatial structure of the inertial range, usually expected to have a wavenumber dependence close to a Kolmogorov-like spectral index of 5/3-5/3. Nevertheless, we can recognize the inertial range form as near $f^{-5/3}$, in essentially all wind categories but the fast solar wind (see below). We also note that toward low frequencies at around $10^{-5}$ to $10^{-4}\,\mathrm{Hz}$, the spectral form gradually transitions to the $1/f$ regime. This extended transition region is consistent with the use of a long data record, along with the established large-variance log-normal distribution of interplanetary magnetic field correlation times (Ruiz et al., 2014; Wang et al., 2024a) and the nature of the superposition principle (Machlup, 1981; Montroll & Shlesinger, 1982) that provides a generic pathway to the generation of $1/f$ signal.
In the fast solar wind ($V_{\mathrm{SW}}\geq 500\,\mathrm{km/s}$), the inertial range spectral index – typically observed at frequencies above around $3\times 10^{-4}\,\mathrm{Hz}$ (if considering the lower bound of the 1 au correlation time distribution at around 1 hour) – deviates from 5/3-5/3 and flattens. This trend is evident across all three magnetic field components. Evidence of spectral flattening in fast wind has been presented by Borovsky (2012) employing ACE intervals categorized by wind speed and wind origin (e.g. coronal-hole, non-coronal-hole, and ejecta based on ionized oxygen ratio). Magnetic field spectral index is systematically less negative in fast wind and in wind of coronal-hole origin. Borovsky (2012) also identifies intervals of sustained shallow spectra reaching a spectral index of -1.34, which tend to have substantially enhanced velocity and magnetic field fluctuations, Alfvén ratios, and outward-to-inward Elsässer ratios. Teodorescu et al. (2015) using measurements at 0.72 au also find spectral flattening in fast wind, and attributes the observed slope tending toward 3/2-3/2 to strong Alfvénicity of the flow and limited turbulence development due to fast wind’s rapid transit. More recently, Sioulas et al. (2023) reports spectral indices near 3/2-3/2 independent of plasma speed closer to the Sun using Parker Solar Probe data.
@@ IV Conclusion and Discussion
A compelling hypothesis is that the observed $1/f$ noise (in the time domain) may originate within the solar dynamo (which admits specific spatial scales). $1/f$-type fluctuations are often found in dynamo simulations and laboratory experiments, in analogy with terrestrial and planetary dynamos (see, e.g., Dmitruk et al., 2014, and references therein). While explicit evidence in solar dynamo is less expected due to its 11-year quasi-periodicity, notable examples exist, such as Polygiannakis et al. (2003) that report a $1/f$ signal in long-term sunspot records, and Kostuchenko & Timashev (1998) that find $1/f$ in solar mean magnetic field, solar irradiance, and solar neutrino capture rate.
One possible explanation is that inverse cascade processes occur in solar dynamo, potentially supported by the presence of pseudo-invariants (see Dmitruk & Matthaeus, 2007, and references therein). Numerical studies show that systems undergoing inverse cascade often develop $1/f$ spectra, through competition with direct cascades (Consolini et al., 2015) or through the generation of timescales much longer than the characteristic bandlimited scales (e.g., the eddy turnover time of the largest structure) in a self-organized manner commonly associated with $1/f$ noise (Ponty et al., 2004; Dmitruk et al., 2011).
The connection of the solar inverse cascade process to signals measured in the time domain is not entirely well understood, but may be connected to self-organized criticality (SOC; Bak et al., 1987, 1988) – a framework describing how systems under small perturbation naturally evolve toward dynamical cluster of minimally stable states with no characteristic spatial and temporal scales. The emergence of a $1/f$ signal in such a framework, possibly including in the dynamo, may be related to interaction across scales (Hwa & Kardar, 1992), anomalous relaxation processes (Weron, 1991), as well as finite size effects (Korzeniowska et al., 2023).
@@
On the other hand, previous work (Isaacs et al., 2015) has shown that long data records, such as those used in the present analysis, tend to exhibit long e-folding times in their inertial range autocorrelations. This behavior likely reflects that the underlying decay is not purely exponential, but better described by a stretched exponential form, consistent with the presence of a broad range of correlation scales. Furthermore, within the framework of self-organized criticality (Bak et al., 1987; Hwa & Kardar, 1992), interactions among these timescales possibly associated with dynamics in the solar dynamo may produce even longer scales (reminiscent of inverse cascade processes), and thereby sustain an extended regime of $1/f$ spectrum.